OpenStudy (anonymous):

is this answer correct? problem: Find the product: (x-6y)(x+6y) my answer is: x^3-4x^2+6x-24=0

OpenStudy (maheshmeghwal9):

no:/

OpenStudy (anonymous):

can you please show me how to get it then?

OpenStudy (anonymous):

No the correct answer is x^2 - 36y^2 Just apply the formula of (a-b)(a+b)=a^2- b^2

OpenStudy (anonymous):

gee i was way off then can i show you another one to see if it is right?

OpenStudy (anonymous):

problem : (4X+4Y)^2 answer : 16x^2+32xy+16y^2

OpenStudy (anonymous):

Use the difference of two squares formula:\[(\color{red} a+\color{red} b)(\color{blue}a-\color{blue}b)=\color{red} a\color{blue}a+\cancel{\color{red} a\color{blue}b}-\cancel{\color{red} a\color{blue}b}+\color{red} b\color{blue}b=a^2-b^2\]

OpenStudy (maheshmeghwal9):

ya:) \[\LARGE{\color{red}{(a+b)(a-b)=a^2-b^2.}}\]So; \[(x-6y)(x+6y)=(x^2)-(36y^2)=x^2-36y^2.\]\[\color{green}{\text{It is ur answer.}}\]

OpenStudy (anonymous):

Use the difference of two squares formula:\[(\color{red} a+\color{red} b)(\color{blue}a-\color{blue}b)=\color{red} a\color{blue}a+\cancel{\color{red} a\color{blue}b}-\cancel{\color{red} a\color{blue}b}+\color{red} b\color{blue}b=a^2-b^2\]

OpenStudy (anonymous):

nope...(x-6y)(x+6y) foiled correctly it would be x^2-36y^2

OpenStudy (maheshmeghwal9):

gt it?

OpenStudy (anonymous):

i was off on that one too. I will go back and re read thank you

OpenStudy (maheshmeghwal9):

k! yw ^_^